In given diatomic gas sound travels at 300m/sec. The root mean square velocity of the molecules of the gas is
a)4m/sec.
b)40m/sec
c)438m/sec
d)43900m/sec

Grade:Select Grade

3 Answers

Aziz Alam IIT Roorkee
8 years ago
v=\sqrt{\gamma P/\rho } velocity of sound in a gas

PM=\rho RT

P/\rho =RT/M

\Rightarrow v=\sqrt{\gamma RT/M}

\Rightarrow RT/M=v^{2}/\gamma

V_{rms}=\sqrt{3RT/M}

\Rightarrow V_{rms}=\sqrt{3v^{2}/\gamma }

Vrms=439.155 m/s

Aziz Alam IIT Roorkee
8 years ago
[v=\sqrt{\gamma P/\rho }] velocity of sound in a gas

PM=\rho RT

P/\rho =RT/M

\Rightarrow v=\sqrt{\gamma RT/M}

\Rightarrow RT/M=v^{2}/\gamma

V_{rms}=\sqrt{3RT/M}

\Rightarrow V_{rms}=\sqrt{3v^{2}/\gamma }

Vrms=439.155 m/s
Karan Shinde
8 years ago
thank you.